OpenStudy (kamille):

If \[\frac{ \sin(a)-\cos(a) }{ \cos(a) }=2\] then \[tg(a) \] is?

hartnn (hartnn):

separate out the numerator

OpenStudy (kamille):

I thought I could: \[\cos(a)(\sin(a)-1)=2\] but I am not sure if it helps?

hartnn (hartnn):